Finding three ordered pairs that satisfy equation. X + Y = 7
I have tried using Algebrator on this but I cannot figure out where 3 ordered pairs come from. (0,7) (7,0) yes but what is the last? (X + y = 7)

Answer by prince_abubu(198)

Practically, all it's saying is what two numbers added together equals 7. The coordinate point (0, 7) does it because 0 + 7 = 7. So does (7, 0) because 7 + 0 = 7.
But so does (1,6) because 1 + 6 = 7. And so does (2,5) because 2 + 5 = 7. There are actually way more than three ordered pairs of numbers that satisfy the equation X + Y = 7.